# Investigation of ArUco Marker Placement for Planar Indoor Localization

### v1

@@ II-A Fiducial markers and ArUco
There exists a large number of fiducial marker based systems. A comprehensive overview and timeline for fiducial markers can be found in [11]. The most popular marker systems are the planar, square, black and white fiducial markers AprilTag [12, 13, 14] and ArUco [9, 15]. A recent comparison between different planar fiducial markers is given in [16], where ArUco shows superior accuracy and speed for robotics applications. In contrast, ArUco gives worse result in [17], but still competitive performance with similar markers (including AprilTag) at low computational cost. The authors of [11] recommend ArUco for general applications due to it’s good software. Besides, ArUco’s software is actively improved, e.g. w.r.t. processing time in ArUco3 [15]. It is also equipped with simultaneous localization and mapping (SLAM) functionality. At first, offline camera localization and marker mapping was proposed in [18]. This was later extended for online processing [19], inclusion of keypoints to the map [20], and utilizing maps built with cameras different from those used during inference [21]. In addition, parts of ArUco are available in OpenCV [10], whose implementation we use in this work. ArUco offers extensive capabilities and easy setup, making it a popular choice for robotics practitioners, and suitable for our study.
@@ III-A Pinhole camera model
For camera pose estimation, this work makes use of the pinhole camera model [31, p. 153ff.], which assumes that an image is generated by rays of light passing through the camera’s pinhole and hitting the sensor plane consisting of photo diodes. We define a 3D point in camera coordinates as $\underline{p}_{w}=[x_{w},y_{w},z_{w}]^{T}\in\mathbb{R}^{3}$. The corresponding homogeneous coordinates are $\underline{\overline{p}}_{w}=[x_{w},y_{w},z_{w},1]^{T}\in\mathbb{R}^{4}$. Its 2D projection in image coordinates $\underline{p}_{i}=[x_{i}/z_{i},y_{i}/z_{i}]^{T}\in\mathbb{R}^{2}$ can be computed from the homogeneous representation $\overline{\underline{p}}_{i}=[x_{i},y_{i},z_{i}]^{T}\in\mathbb{R}^{3}$. The projection of corresponding points in world coordinates $\underline{\overline{p}}_{w}$ to image coordinates $\underline{\overline{p}}_{i}$ is given by
@@ III-C Camera calibration
The intrinsic matrix $\mathbf{K}$ is determined offline through camera calibration. During localization, $\mathbf{K}$ is passed to the OpenCV PnP solver to obtain the extrinsic parameters. Camera calibration requires at least ten images from different perspectives of a known calibration pattern. The images are then used to first estimate the homography for each view and then the intrinsic parameters [32, p. 665f.]. Commonly used calibration patterns include a chessboard,
ArUco or ChArUco patterns [32, p. 652ff.]. During this work, the chessboard pattern was used, as it led to the best pose estimates. Calibration with a liquid-crystal display (LCD) instead of a printed version, as suggested in [33], was examined and did not bring improvements in localization.
@@ III-D Marker detection
To solve the PnP problem, correspondences between points in the image plane and in the world coordinate system are required. While the world points are measured during marker installation, the image points are detected with the provided ArUco software in OpenCV. For corner refinement, we apply the CORNER REFINE CONTOUR method [10] based on line fitting of detected contour points of the fiducial marker. After using the outer border of the markers for initial detection, and the inner pattern for determination of marker IDs and rotations, the four corners of the quadratic markers are returned and can be used as detected 2D image plane points $\underline{\hat{p}}_{ik}$ in the PnP solver.
@@ VI-A Influence of camera-marker distance, marker orientation, and distance from image center
The two different camera-marker distances in configuration I show the expected increase in localization error with larger camera-marker distances, as evident when comparing Table II with Table III, and Fig. 3(a) with Fig. 3(b). Fig. 3(b) misses a few outliers. It can be observed in both figures and tables that inner markers perform worse than the outer markers. In the experiments conducted for Fig. 3(b), 50% of the markers on the inner circle show errors greater than $20\,\text{cm}$ (with many outliers not included in the figure), whereas for the outer circle it’s only 31.25%. This suggests that markers placed further from the image center are more beneficial for localization, likely because inner markers are closer to the problematic frontal marker observation. In additional experiments with two-marker placements, distributing them over the image space and placing them far from the center also gave the best results, as the markers cover a larger geometric aperture.
@@ VI-C Kalman filter tracking
For the experiments with our proposed Kalman filter in configuration II, the markers are placed on the ceiling and the camera is moved (by hand) on the floor as shown in Fig. 2. The resulting ground truth is visualized in Fig. 6 together with the estimated camera positions from camera pose estimation with measured fiducial markers and the posteriors $\underline{x}_{t}$ from (14) for the different measurement variance models. The max, mean and median variances are too large and cause notable deviations from the ground truth. The min variances lead to accurate AMR trajectories if enough markers are visible (bottom with $\leq 5$ and right with all visible markers). However, when less markers are visible (left $\leq 3$ and top $\leq 1$), the min variances cause nonsmooth trajectories, as visible in the zoomed-in region, because the Kalman filter becomes overconfident in the measurements. Our proposed adaptive measurement variances give high accuracy and smooth AMR motion in all regions and overall best performance.
Figure 6: Kalman filter tracking results for the rectangular trajectory. Blue dots are the estimated camera position inputs to the Kalman filter.

### v2

@@ Abstract
Note: This is the accepted version of the paper. The final version will be published in the Proceedings of the IEEE 2025 22nd International Conference on Advanced Robotics (ICAR).
footnotetext: © 2025 Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works
@@ II-A Fiducial markers and ArUco
There exists a large number of fiducial marker based systems. A comprehensive overview and timeline for fiducial markers can be found in [11]. The most popular marker systems are the planar, square, black and white fiducial markers AprilTag [12, 13, 14] and ArUco [9, 15]. A recent comparison between different planar fiducial markers is given in [16], where ArUco shows superior accuracy and speed for robotics applications. In contrast, ArUco gives worse result in [17], but still competitive performance with similar markers (including AprilTag) at low computational cost. The authors of [11] recommend ArUco for general applications due to its good software. Besides, ArUco’s software is actively improved, e.g. w.r.t. processing time in ArUco3 [15]. It is also equipped with simultaneous localization and mapping (SLAM) functionality. At first, offline camera localization and marker mapping was proposed in [18]. This was later extended for online processing [19], inclusion of keypoints to the map [20], and utilizing maps built with cameras different from those used during inference [21]. In addition, parts of ArUco are available in OpenCV [10], whose implementation we use in this work. ArUco offers extensive capabilities and easy setup, making it a popular choice for robotics practitioners, and suitable for our study.
@@ III-A Pinhole camera model
For camera pose estimation, this work makes use of the pinhole camera model [31, p. 153ff.], which assumes that an image is generated by rays of light passing through the camera’s pinhole and hitting the sensor plane consisting of photo diodes. We define a 3D point in world coordinates as $\underline{p}_{w}=[x_{w},y_{w},z_{w}]^{T}\in\mathbb{R}^{3}$. The corresponding homogeneous coordinates are $\underline{\overline{p}}_{w}=[x_{w},y_{w},z_{w},1]^{T}\in\mathbb{R}^{4}$. Its 2D projection in image coordinates $\underline{p}_{i}=[x_{i}/z_{i},y_{i}/z_{i}]^{T}\in\mathbb{R}^{2}$ can be computed from the homogeneous representation $\overline{\underline{p}}_{i}=[x_{i},y_{i},z_{i}]^{T}\in\mathbb{R}^{3}$. The projection of corresponding points in world coordinates $\underline{\overline{p}}_{w}$ to image coordinates $\underline{\overline{p}}_{i}$ is given by
@@ III-C Camera calibration
The intrinsic matrix $\mathbf{K}$ is determined offline through camera calibration. During localization, $\mathbf{K}$ is passed to the OpenCV PnP solver to obtain the extrinsic parameters. Camera calibration requires at least ten images from different perspectives of a known calibration pattern. The images are then used to first estimate the homography for each view and then the intrinsic parameters [32, p. 665f.]. Commonly used calibration patterns include a chessboard, ArUco or ChArUco patterns [32, p. 652ff.]. During this work, the chessboard pattern was used, as it led to the best pose estimates. Calibration with a liquid-crystal display (LCD) instead of a printed version, as suggested in [33], was examined and did not bring improvements in localization.
@@ III-D Marker detection
To solve the PnP problem, correspondences between points in the image plane and in the world coordinate system are required. While the world points are measured during marker installation, the image points are detected with the provided ArUco software in OpenCV. After detecting the outer border of the marker, the inner pattern is extracted for determination of marker ID and rotation and the four corners of the quadratic marker are returned. For corner refinement, we apply the CORNER REFINE CONTOUR method [10] based on line fitting of detected contour points of the fiducial marker. These four image points are then used as detected 2D image plane points $\underline{\hat{p}}_{ik}$ in the PnP solver.
@@ VI-A Influence of camera-marker distance, marker orientation, and distance from image center
The two different camera-marker distances in configuration I show the expected increase in localization error with larger camera-marker distances, as evident when comparing Table II with Table III, and Fig. 3(a) with Fig. 3(b). Fig. 3(b) misses outliers. It can be observed in both figures and tables that inner markers perform worse than the outer markers. In the experiments conducted for Fig. 3(b), 50% of the markers on the inner circle show errors greater than $20\,\text{cm}$ (with many outliers not included in the figure), whereas for the outer circle it’s only 31.25%. This suggests that markers placed further from the image center are more beneficial for localization, likely because inner markers are closer to the problematic frontal marker observation. In additional experiments with two-marker placements, distributing them over the image space and placing them far from the center also gave the best results, as the markers cover a larger geometric aperture.
@@ VI-C Kalman filter tracking
For the experiments with our proposed Kalman filter in configuration II, the markers are placed on the ceiling and the camera is moved (by hand) on the floor as shown in Fig. 2. The resulting ground truth is visualized in Fig. 6 together with the estimated camera positions from camera pose estimation with measured fiducial markers and the posteriors $\underline{x}_{t}$ from (14) for the different measurement variance models. The max, mean and median variances are too large and cause notable deviations from the ground truth, as seen in Fig. 6(a). The min variances lead to accurate AMR trajectories if enough markers are visible (bottom with $\leq 5$ and right with all visible markers). However, when less markers are visible (left $\leq 3$ and top $\leq 1$), the min variances cause nonsmooth trajectories as shown in Fig. 6(b). There, the Kalman filter becomes overconfident in the less reliable measurements. Our proposed adaptive measurement variances give high accuracy and smooth AMR motion in all regions and overall best performance.
((a)) Best tracking performance with min and adaptive (Ada.) variance models.
((b)) Adaptive variance model is more robust than min variance model.
Figure 6: Kalman filter tracking results for a) the rectangular trajectory and b) the start region with single-marker positioning. Black dots are the estimated camera position inputs to the Kalman filter. The gray rectangles denote the hand-measured ground truth with 5 cm confidence interval.